OVHCloud Bare Metal Cloud Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#7364 — VPS robots
Incident Report for Bare Metal Cloud
Resolved
We notice perturbations on VPS management robots.
All robots are stopped the time of diagnosis.

This affects all tasks related to VPS as well as graphs of statistics in the managerv5.

Update(s):

Date: 2012-09-25 20:41:18 UTC
The service is UP.

Date: 2012-09-25 20:39:04 UTC
The robots are now relaunched, we are fixing the last tasks in error.
The infrastructure is operational again.

Date: 2012-09-25 20:38:11 UTC
The clusters are working again.
We are going to reactivate the robots progressively.

Date: 2012-09-25 20:37:36 UTC
The hosts are reconnected successfully.
We are going to reactivate all the functionalities of HA on the clusters.

Date: 2012-09-25 20:36:41 UTC
We reconnect the hosts to vCenter, the statistics should start to go back progressively for the whole VPS.
We didn't found particular problems for the time being.

Date: 2012-09-25 20:34:28 UTC
The update started, the vCenter is updated and started but the hosts are still not completely added.
We keep progressing.

The VPS are not absolutely impacted, only the management of the infrastructur is momentarily suspended.

Date: 2012-09-25 20:32:01 UTC
We still have problems to restart the vCenter VPS, the robots will remain interrupted for the moment.
We are going to update the vCenter in order to manage better the size of the infra and take advantage of the last updates.
Posted Sep 25, 2012 - 09:36 UTC